# Collective order boundedness of sets of operators between ordered vector spaces

### v1

@@ Abstract
It is proved that: each collectively order continuous set of operators from an Archimedean ordered vector space to an ordered vector space is collectively order bounded; each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded; collectively r-continuous sets agree with collectively order bounded ones when the space in the range has a generating cone. Several applications to operator semigroups are given.
Keywords: ordered vector space, collectively qualified set of operators, operator semigroup
@@ 1 Introduction
Collectively compact sets of operators between normed spaces were introduced by P. M. Anselone and T. W. Palmer [5]. Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 7, 8, 9]. The present paper generalizes results of [7] to the ordered vector space setting. The results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator between Archimedean ordered vector spaces is order bounded; 2) each order to norm continuous operator from an ordered Banach space with a closed generating cone to a normed space is continuous.
@@
    relative uniform convergent to $x\in X$ (briefly, r-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
@@
    collective r-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
@@
    ${\cal T}$ is collectively r-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
@@
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an OVS to an Archimedean OVS is collectively order bounded. Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets. Finally, we prove Theorem 2.3 saying that under rather mild conditions, every collectively order to norm bounded set of operators is norm bounded. Section 3 is devoted to applications of obtained results to one-parameter operator semigroups on OVSs.
@@ 2 Main results
We are going to prove the following results which extend the results obtained recently in [7] in the vector lattice setting.
@@
    Every collectively order continuous set of operators from an Archimedean OVS to an OVS is collectively order bounded.
@@
We begin with an extension of the first part of [7, Theorem 2.2].
@@
Let $X$ and $Y$ be OVSs, with an Archimedean $X$. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
@@
In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] in the following way.
@@
Each order continuous operator from an Archimedean OVS to an OVS is order bounded.
The Archimedean assumption is essential in Corollary 2.1.
@@
Next, we extend the last part of [7, Theorem 2.2] to OVSs.
@@
Let $X$ and $Y$ be OVSs then $\text{\bf L}_{rc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$. If additionally $Y=Y_{+}-Y_{+}$ then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
@@
We do not know whether the assumption $Y=Y_{+}-Y_{+}$ in Theorem 2.2 is essential for $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
@@
Order bounded operators from an OVS to an OVS with generating positive cone agree with r-continuous operators.
@@
Let $X$ be an OVS and $Y$ be an OBS such that $Y_{+}$ is closed, generating, and normal. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
@@
Let $X$ be an ordered Banach space with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
@@
The condition $X_{+}-X_{+}=X$ is essential in Theorem 2.3.
@@
3 Applications to commutative operator semigroups
We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case). More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [10, p.75]).
@@ 3 Applications to commutative operator semigroups
Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ containing a (weakly) compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
@@
Combining Propositions 3.1 and 2.1 we obtain.
@@
Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [10, Theorem 1.5]). Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
@@
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a discrete operator semigroup on an order continuous BL $E$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.

### v2

@@ Abstract
It is proved that: each collectively order continuous set of operators between Archimedean ordered vector spaces with generating cones is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded. Several applications to commutative operator semigroups on ordered vector spaces are given.
Keywords: ordered vector space, collectively qualified set of operators, commutative operator semigroup
@@ 1 Introduction
Collectively compact sets of operators between normed spaces were introduced by P. M. Anselone and T. W. Palmer [5]. Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 7, 8, 9]. The present paper generalizes results of [1, 7, 15] to the ordered vector space setting. The results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator between Archimedean ordered vector spaces is order bounded; 2) each order to norm continuous operator from an ordered Banach space with a closed generating cone to a normed space is continuous.
@@
    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
@@
    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
@@
    ${\cal T}$ is collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
@@
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an OVS to an Archimedean OVS is collectively order bounded. Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets. Finally, we prove Theorem 2.3 saying that under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded. It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded. Section 3 is devoted to applications of obtained results to commutative operator semigroups on OVSs.
@@ 2 Main results
We are going to prove the following results which generalize the results obtained recently in [7, 15] in the vector lattice setting to the OVS setting.
@@
    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS with a generating cone is collectively order bounded.
@@
We begin with generalization of the first part of [7, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
@@
Let $X$ and $Y$ be OVSs with generating cones and suppose $X$ is Archimedean. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
@@
In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] as follows.
@@
Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $Y$ with generating cone is order bounded.
The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Corollary 2.1 even in the VL setting.
@@
Next, we generalize [15, Theorem 5.1] and [7, Theorem 2.2] to OVSs.
@@
Let $X$ and $Y$ be OVSs with generating cones then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
@@
Order bounded operators between OVSs with generating cones agree with r-continuous operators.
@@
Let $X$ be an OVS and $Y$ be an OBS with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
@@
Let $X$ be an OBS with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
@@
It is worth noting that even in the case of a single operator, Theorem 2.3 generalizes the well known fact that every order bounded operator between BLs is bounded. The condition $X_{+}-X_{+}=X$ is essential in Theorem 2.3.
@@
3 Some applications to operator semigroups
We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case). More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [13, p.75]).
@@ 3 Some applications to operator semigroups
Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
@@
Combining Propositions 3.1 and 2.1, and keeping in mind that every OBS with a generating normal cone is Archimedean, we obtain.
@@
Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [13, Theorem 1.5]). Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
@@
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a one-parameter discrete operator semigroup on an order continuous BL $E$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on OVSs.
The next two lemmas follow directly from [7, Theorems 2.1 and 2.2] and from Theorems 2.1, 2.2, and 2.3.
Lemma 3.1.
Every collectively order to norm bounded set of operators on an (OBS) with a closed generating cone is norm bounded.
Lemma 3.2.
Let $X$ be and OVS with generating cone. Then $\text{\bf L}_{rc}(X)=\text{\bf L}_{ob}(X)$. If $X$ is additionally Archimedean then $\text{\bf L}_{oc}(X)\subseteq\text{\bf L}_{ob}(X)$.
Next definition localizes several notions concerning semigrops.
Definition 3.1.
A semigroup $(T_{s})_{s\geq 0}$ on an OVS $X$ is
  1. $a)$
    locally order bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M$.
  2. $b)$
    locally o-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M$.
  3. $c)$
    locally ru-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{rc}(X)$ for every $M$.
A semigroup $(T_{s})_{s\geq 0}$ on a normed OVS $X$ is
  1. $e)$
    locally order to norm bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$
    for every $M$.
In several important cases, in each of the four items of Definition 3.1 the condition “for every $M$” can be replaced by “for some $M$”.
Proposition 3.3.
Let $(T_{s})_{s\geq 0}$ be a semigroup on an OVS $X$ such that $X_{+}-X_{+}=X$.
  1. $i)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order bounded.
  2. $ii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{rc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally ru-continuous.
  3. $iii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order continuous.
  4. $iv)$
    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order to norm bounded.
Proof.
$i)$ Let $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Let $n\varepsilon\geq M$, and take arbitrary $a\leq b$ in $X$. It follows from $X_{+}-X_{+}=X$ that $[a,b]\subseteq[-a_{0},a_{0}]$ for some $a_{0}\in X_{+}$. Since $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ and $X_{+}$ is generating, $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-a_{1},a_{1}]$ for some $a_{1}\in X$. Similarly, for some $a_{2}\in X$, we have $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}]$ and
$\bigcup\limits_{\varepsilon\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]=T_{\varepsilon}\left(\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\right)\subseteq T_{\varepsilon}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}].$
It follows $\bigcup\limits_{0\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-(a_{1}+a_{2}),a_{1}+a_{2}]$. Repeating n-times the argument, we obtain $a_{1},a_{2},...,a_{n}\in X$ with
$\bigcup\limits_{0\leq s\leq M}T_{s}[a,b]\subseteq\bigcup\limits_{0\leq s\leq n\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq\left[-\sum_{i=1}^{n}a_{i},\sum_{i=1}^{n}a_{i}\right].$ (1)1( 1 )
Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order bounded.
$ii)$ By Lemma 3.2, it follows from $i)$.
$iii)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, and fix $M\geq 0$. Let $n\varepsilon\geq M$ and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Pick a net $g^{(1)}_{{\beta_{(1)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(1)}$ there is an $\alpha_{\beta_{(1)}}$ with ±$\pm T_{s}x_{\alpha}\leq g^{(1)}_{{\beta_{(1)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(1)}}$. Since $T_{\varepsilon}$ is order continuous, $T_{\varepsilon}x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. It follows from $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ that there exists a net $g^{(2)}_{{\beta_{(2)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(2)}$ there is $\alpha_{\beta_{(2)}}$ with ±$\pm T_{s}T_{\varepsilon}x_{\alpha}\leq g^{(2)}_{{\beta_{(2)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(2)}}$. Repeating n-times, we obtain nets
$g^{(1)}_{{\beta_{(1)}}}\downarrow 0,\ g^{(2)}_{{\beta_{(2)}}}\downarrow 0,\ ...\ g^{(n)}_{{\beta_{(n)}}}\downarrow 0$
so that, for each $\beta_{(i)}$, $i=1,2,...,n$, there is $\alpha_{\beta_{(i)}}$ satisfying
±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ \alpha\geq\alpha_{\beta_{(i)}}).$
Consider the net $g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}=\sum_{i=1}^{n}g^{(i)}_{{\beta_{(i)}}}\downarrow 0$. Take an
$\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\geq\alpha_{\beta_{(1)}},\alpha_{\beta_{(2)}},....,\alpha_{\beta_{(n)}}.$
Then, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ i=1,2,...,n).$
Therefore, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
±$\pm T_{s}x_{\alpha}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq n\varepsilon).$ (2)2( 2 )
Since $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ is arbitrary, (2) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq n\varepsilon}\in\text{\bf L}_{oc}(X)$, and hence $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order continuous.
$iv)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Let $n\varepsilon\geq M$ and take some $a\in X_{+}$. By Lemma 3.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$. Then
$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C\|a\|.$ (3)3( 3 )
Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order to norm bounded. ∎
In recent works [10, 11, 12] relatively uniformly continuous semigroups on VLs were deeply studied. The use of collectively qualified semigroups is to rough for getting progress in this direction. We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.